OpenStudy (anonymous):

Find the slope of the line that passes through the points (0, 0) and (-2, -3). 2/3 3/2 -3/2

OpenStudy (raden):

use the formula : m = (y2-y1)/(x2-x1) with m is slope

OpenStudy (anonymous):

3/2 the equation of the line is y = mx + b . Replace the the point (0,0) to the equation you have b = 0.Then you have the equation is y = mx. again replace the point (-2.-3) to the equation you have m = y / x it is m = -3/ -2 = 3/2